OpenStudy (anonymous):

Solve (29/x) - (2/x) < or equal to 3

OpenStudy (anonymous):

what does \[\frac{ 29 }{ x } - \frac{ 2 }{ x } =?\]

OpenStudy (anonymous):

27/x

OpenStudy (anonymous):

so \[\frac{ 27 }{ x } \le 3 \]

OpenStudy (anonymous):

how do you get x out of the denominator?

OpenStudy (anonymous):

multiply both sides by x then divide both sides by 3

OpenStudy (anonymous):

yep